# On Geometry and Topology of Hessian Manifolds

### v3

@@ Abstract
Key topological constraints are derived, including the vanishing of the Euler characteristic for any compact Hessian manifold and infinite torsion-freeness of its fundamental group. We demonstrate that compact Hessian manifolds with an abelian fundamental group are homeomorphic to the flat torus. The theory of flat line bundles is developed to introduce the canonical and obstruction bundles, which are used to prove that a curved compact orientable Hessian manifold must be a mapping torus. Curvature properties are then discussed. As an application, we prove that all compact Hessian-Einstein manifolds are flat.
A Hessian manifold $((M,\nabla),g)$ is said to be of Koszul type, if there exists a $1$-form $\eta\in\Omega^{1}(M)$ such that $g=\nabla\eta$. We place emphasis on Hessian manifolds of Koszul type, establishing an equivalence between their existence and that of positive flat line bundles, space-like Lagrangian embeddings, and hyperbolic affine structures. A duality with radiant manifolds, analogous to the classical Legendre transformation, is explored. We also construct a non-trivial family of compact Hessian manifolds of Koszul type with negative scalar curvature, and classify all complete Hessian surfaces.
@@ 1.1 Introduction and Background
Having established this broad landscape, we switch from these foundational constraints to a more refined analysis by developing the theory of flat line bundles as our primary analytical tool. We introduce this as a novel framework for studying Hessian geometry and define two crucial new objects within it: the canonical line bundle, an intrinsic invariant of the affine structure, and the obstruction line bundle. The latter is constructed to precisely measure the failure of locally defined Hessian potentials to patch together into a global function; we show its gauge theoretic triviality is equivalent to the existence of a global potential function. As a direct consequence, we also establish that any orientable compact Hessian manifold with vanishing first Betti number must be flat (Theorem2.40).
@@
A central achievement of our work is to reveal that Hessian metrics of Koszul-type are part of a much larger network of connections than previously understood. Whilst the work of H.Shima [15] established an important equivalence between the existence of a Koszul type Hessian metric and that of a hyperbolic affine structure, we establish a new, powerful list of equivalences, demonstrating that these conditions are also entirely equivalent to two other seemingly disparate properties: the existence of a positive flat line bundle (Corollary2.49) and admitting a space-like Lagrangian embedding (Corollary3.30) into its own cotangent bundle, equipped with the Riemann extension metric first studied by Patterson and Walker in 1952 [16], thereby building a new bridge between affine differential geometry and para-Kähler geometry.
@@
The synthesis of these tools now culminates in powerful structural theorems. By applying D.Tischler’s classical 1970 theorem on fibrations [19] to the first Koszul form, we prove that any curved compact orientable Hessian manifold must be a mapping torus (Theorem2.43), revealing a remarkably rigid underlying structure, particularly in low dimensions (Corollary2.44). This result is further strengthened by a de Rham-type splitting theorem, which shows that the universal cover of a compact Hessian manifold isometrically splits as a product of another contractible Riemannian manifold and the real line (Proposition3.56). As a important application, we then prove that all compact Hessian-Einstein manifolds are flat (Theorem3.60). An interesting geometric uncertainty principle is also established on Hessian manifolds, revealing links to quantum mechanics.
After obtaining a number of theorems, we recall that constructing explicit examples of non-trivial curved compact Hessian manifolds has been a significant challenge, with most known examples being flat Bieberbach manifolds. We therefore address this by explicitly constructing a non-trivial family of compact radiant Hessian manifolds of Koszul type that possess negative scalar curvature 1-1, suggesting that our theory is not vacuous.
Finally, as one of the highlights of this article, we adopt our general theory to low dimensions. Applying classical algebraic topology techniques and the solvability of Heisenberg group $H_{3}(\mathbb{R})$, we provide the first complete topological classification of all complete Hessian surfaces (Theorem4.5). We also derive severe constraints on the intersection form for Hessian $4$-manifolds, and then prove that any Einstein metric on a closed Hessian $4$-manifold must be flat (Proposition4.21). Last but not least, a novel characterization for potentials of $2$-dimensional flat metrics is established, showing that a pseudo-Riemannian metric
@@
This paper is organized as follows: Section1 reviews fundamental definitions and introduces a warped product construction as first examples. Section2 suggests key topological obstructions, concerning the vanishing of Euler characteristic and the infinitude of fundamental group. We then develop the theory of flat line bundles, proving that curved compact orientable Hessian manifolds must be mapping tori. Section3 establishes the generalised Legendre transform, and a central equivalence between Hessian metrics of Koszul type and several other geometric structures, including positive flat line bundles and space-like Lagrangian embeddings. Curvature properties and link to complex analysis are also discussed. Finally, Section4 applies these results to low dimensions, yielding a complete topological classification of Hessian surfaces and deriving strong constraints for Hessian 4-manifolds. An appendix is included to support our classification of compact Hessian manifolds with Abelian fundamental groups.
@@ 2.2 Fundamental Group and Euler Number
Combine Theorem2.13 and Theorem2.15&2.16, we immediately get the following result, confirming Conjecture1.6 for Hessian manifolds.
@@ 2.3 Flat Line Bundles
2.3.1 The Picard group
In this section, we study Hessian manifolds using the concept of flat line bundles. We also define the Picard group, and use it to prove that a curved compact orientable Hessian manifold has positive first Betti number.
@@ 2.3.1 The Picard group
Let $\mathscr{L}\rightarrow M$ be a $\mathbb{R}$-line bundle over a differentiable manifold $M$. By the local triviality of $\mathscr{L}\rightarrow M$, there exists an open covering $\{U_{\alpha}\}_{\alpha\in I}$ of $M$ and for each $\alpha\in I$ a local section $s_{\alpha}$ of $\mathscr{L}\rightarrow M$ defined on $U_{\alpha}$, such that $s_{\alpha}$ is non-vanishing for all $\alpha\in I$. The collection $\{s_{\alpha}\}_{\alpha\in I}$ is traditionally termed a local frame field of $\mathscr{L}\rightarrow M$.
@@
Let $M$ be a differentiable manifold and let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$. Then $\mathscr{L}\rightarrow M$ admits a local frame field $\{s_{\alpha}\}_{\alpha\in I}$ such that $Ds_{\alpha}=0$ for all $\alpha\in I$. Moreover, for each $\alpha,\beta\in I$ there exists a constant $g_{\alpha\beta}\in\mathbb{R}^{\times}$ such that
$s_{\beta}(p)=s_{\alpha}(p)\cdot g_{\alpha\beta}$
for all $p\in U_{\alpha}\cap U_{\beta}$, where $U_{\alpha}$ here is the domain of definition of $s_{\alpha}$ and $U_{\beta}$ is the domain of definition of $s_{\beta}$.
@@
Let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over a differentiable manifold $M$. A local frame field $\{s_{\alpha}\}_{\alpha\in I}$ of $\mathscr{L}\rightarrow M$ is termed an affine local frame field of $(\mathscr{L},D)$, if $Ds_{\alpha}=0$ for all $\alpha\in I$.
@@
Two flat $\mathbb{R}$-line bundles $(\mathscr{L}_{1},D_{1})$ and $(\mathscr{L}_{2},D_{2})$ over $M$ are said to be gauge equivalent, if there exists a $\mathbb{R}$-vector bundle isomorphism $\phi\colon\mathscr{L}_{1}\rightarrow\mathscr{L}_{2}$ such that $D_{1}-\phi^{*}D_{2}\in\Omega^{1}(M)$ is an exact 1-form on $M$.
Theorem 2.27.
Let $M$ be a differentiable manifold. Up to gauge equivalence, flat $\mathbb{R}$-line bundles over $M$ are classified by $\operatorname{Hom}(\pi_{1}(M),\mathbb{R}^{\times})$, or equivalently, by the additive group $\mathrm{H}^{1}(M;\mathbb{R})\oplus\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})$ if $M$ is aspherical.
@@
Notice that by the universal coefficients and Hurewicz’s theorem,
$\mathrm{H}^{1}(M;\mathbb{R})\oplus\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})\cong\operatorname{Hom}(\pi_{1}(M),\mathbb{R})\oplus\operatorname{Hom}(\pi_{1}(M),\mathbb{Z}/2\mathbb{Z})=\operatorname{Hom}(\pi_{1}(M),\mathbb{R}^{\times})$
where $\pi_{1}(M)$ is the fundamental group of $M$.
Let $\rho\colon\pi_{1}(M)\rightarrow\mathbb{R}^{\times}$ be a representation. Consider the associated bundle $\mathscr{L}:=\tilde{M}\times_{\rho}\mathbb{R}$ where $\pi\colon\tilde{M}\rightarrow M$ is a universal cover, and each element $g$ of $\pi_{1}(M)$ acts diagonally by
$g\cdot(p,x):=(g(p),\rho(g)x)$
for every $p\in\tilde{M}$ and $x\in\mathbb{R}$. By construction, the connection $d$ on the trivial bundle is $\pi_{1}(M)$-equivariant, and hence descends to a flat connection on $\mathscr{L}$.
Conversely, for any flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ over $M$, since the curvature of $D$ vanishes identically, the holonomy representation sending a homotopy class represented by a smooth loop $\gamma$ in $M$ to the parallel transport along $\gamma$ is a well-defined group representation $\rho\colon\pi_{1}(M)\rightarrow\operatorname{GL}_{1}(\mathbb{R})=\mathbb{R}^{\times}$.
Moreover, if $D$ be a flat connection on $\mathscr{L}$ and $f\in C^{\infty}(M)$, then the curvature 2-form of $D+df$ is
$ddf+df\wedge df=0$
and by the gradient theorem the holonomy of $D$ coincides with that of $D+df$. ∎
@@
This is essentially a reformulation of Theorem2.27. ∎
Definition 2.29.
Let $M$ be a differentiable manifold. The group of gauge equivalence classes of flat $\mathbb{R}$-line bundles over $M$ is termed the Picard group of $M$.
Recall that a line bundle is trivial if and only if it admits a nowhere vanishing global section. The following proposition gives a refinement in the case of flat line bundles.
@@
2.3.2 The canonical line bundle

### v4

@@ Abstract
Key topological constraints are derived, including the vanishing of the Euler characteristic for any compact Hessian manifold and infinite torsion-freeness of its fundamental group. We demonstrate that compact Hessian manifolds with an abelian fundamental group are homeomorphic to the flat torus. The theory of flat line bundles is developed to introduce the Picard group and the canonical class, which are used to prove that a curved compact orientable Hessian manifold must be a mapping torus. Curvature properties are then discussed. As an application, we prove that all compact Hessian-Einstein manifolds are flat.
A Hessian manifold $((M,\nabla),g)$ is said to be of Koszul type, if there exists a $1$-form $\eta\in\Omega^{1}(M)$ such that $g=\nabla\eta$. We place emphasis on Hessian manifolds of Koszul type, establishing an equivalence between their existence and that of positive flat line bundles, space-like Lagrangian embeddings, and hyperbolic affine structures. A duality with radiant manifolds, analogous to the classical Legendre transformation, is explored. We also develop a geometric uncertainty principle on Hessian manifolds of Koszul type, and classify all complete Hessian surfaces and closed Hessian $3$-folds.
@@ 1.1 Introduction and Background
Having established this broad landscape, we switch from these foundational constraints to a more refined analysis by developing the theory of flat line bundles as our primary analytical tool. We introduce this as a novel framework for studying Hessian geometry and define two crucial new objects within it: the canonical class, an intrinsic invariant of the affine structure, and the locally constant Picard group. The latter classifies all flat $\mathbb{R}$-line bundles over a differentiable manifold up to gauge equivalence. As a direct consequence, we also establish that any orientable compact Hessian manifold with vanishing first Betti number must be flat (Theorem 2.45).
@@
A central achievement of our work is to reveal that Hessian metrics of Koszul type are part of a much larger network of connections than previously understood. Whilst the work of H.Shima [15] established an important equivalence between the existence of a Koszul-type Hessian metric and that of a hyperbolic affine structure, we establish a new, powerful list of equivalences, demonstrating that these conditions are also entirely equivalent to two other seemingly disparate properties: the existence of a positive flat line bundle (Proposition 2.50) and admitting a space-like Lagrangian embedding (Corollary 3.32) into its own cotangent bundle, equipped with the Riemann extension metric first studied by Patterson and Walker in 1952 [16], thereby building a new bridge between affine differential geometry and para-Kähler geometry.
@@
The synthesis of these tools now culminates in powerful structural theorems. By applying D.Tischler’s classical 1970 theorem on fibrations [19] to the first Koszul form, we prove that any curved compact orientable Hessian manifold must be a mapping torus (Theorem 2.48), revealing a remarkably rigid underlying structure, particularly in low dimensions (Corollary 5.15). This result is further strengthened by a de Rham-type splitting theorem, which shows that the universal cover of a compact Hessian manifold isometrically splits as a product of another contractible Riemannian manifold and the real line (Proposition 4.8). Using this splitting structure and tools from geometric analysis, we are able to prove that a compact Hessian manifold cannot support a metric of negative curvature(Theorem 4.10). As yet another important application, we then prove that all compact Hessian-Einstein manifolds are flat (Theorem 4.13). An interesting geometric uncertainty principle is also established on Hessian manifolds, revealing links to quantum mechanics.
After obtaining a number of theorems, we recall that constructing explicit examples of non-trivial curved compact Hessian manifolds has been a significant challenge, with most known examples being flat Bieberbach manifolds. We therefore address this by explicitly constructing a non-trivial family of compact radiant Hessian manifolds of Koszul type that possess negative scalar curvature 1-1, suggesting that our theory is not vacuous. Furthermore, this careful investigation in low dimension finally results in the first complete topological classification of closed orientable Hessian 3-manifolds(Theorem 5.15), namely, they are always mapping tori of periodic mapping classes.
Finally, as one of the highlights of this article, we adopt our general theory to low dimensions. Applying classical algebraic topology techniques and using the solvability of Heisenberg group $H_{3}(\mathbb{R})$, we provide the first complete topological classification of all complete Hessian surfaces (Theorem 5.11). We also derive severe constraints on the intersection form for Hessian $4$-manifolds, and then prove that any Einstein metric on a closed Hessian $4$-manifold must be flat (Proposition 5.22). Last but not least, a novel characterization for potentials of $2$-dimensional flat metrics is established, showing that a pseudo-Riemannian metric
@@
This paper is organized as follows: Section 1 reviews fundamental definitions and introduces a warped product construction as first examples. Section 2 establishes key topological obstructions, such as the vanishing of the Euler characteristic and the infinitude of the fundamental group. We then develop the theory of flat line bundles, proving that curved compact orientable Hessian manifolds must be mapping tori. Section 3 explores the interaction with other geometric fields, establishing a central equivalence between Hessian metrics of Koszul type and several other structures, including positive line bundles and space-like Lagrangian embeddings. A duality via a generalized Legendre transform and a link to complex analysis are also discussed. Section 4 delves into the geometric analysis of Hessian metrics, studying their curvature tensors and proving that a compact Hessian manifold cannot support a metric of negative curvature, and that all compact Hessian-Einstein manifolds are flat. Finally, Section 5 applies these results to low dimensions, yielding a complete topological classification of complete Hessian surfaces and closed orientable Hessian 3-folds, and deriving strong constraints for Hessian 4-manifolds. An appendix on homotopy tori is included to support our classification of compact Hessian manifolds with Abelian fundamental groups, while another appendix is devoted to the geometric analysis techniques used in our study of curvature properties.
@@ 2.2 Fundamental Group and Euler Number
Combine Theorem 2.13 and Theorem 2.152.16, we immediately get the following result, confirming Conjecture 1.6 for Hessian manifolds.
@@ 2.3 Flat Line Bundles
2.3.1 Locally Constant Picard Group
In this section, we study Hessian manifolds using the concept of flat line bundles. We also define and study the locally constant Picard group of a differentiable manifold, and use it to prove that a curved compact orientable Hessian manifold has positive first Betti number.
@@ 2.3.1 Locally Constant Picard Group
To fix the terminology, we recall the definition of a local frame field, and that of its gluing 1-cocycles.
Let $\mathscr{L}\rightarrow M$ be a $\mathbb{R}$-line bundle over a differentiable manifold $M$. By the local triviality of $\mathscr{L}\rightarrow M$, there exists an open covering $\{U_{\alpha}\}_{\alpha\in I}$ of $M$ and for each $\alpha\in I$ a local section $s_{\alpha}$ of $\mathscr{L}\rightarrow M$ defined on $U_{\alpha}$, such that $s_{\alpha}$ is non-vanishing for all $\alpha\in I$. The collection $\{s_{\alpha}\}_{\alpha\in I}$ is traditionally termed a local frame field of $\mathscr{L}\rightarrow M$. For each $\alpha,\beta\in I$, there exists a unique nowhere vanishing smooth function $g_{\alpha\beta}\in C^{\infty}(U_{\alpha}\cap U_{\beta})$ such that
$s_{\beta}(p)=s_{\alpha}(p)\cdot g_{\alpha\beta}(p)$
for all $p\in U_{\alpha}\cap U_{\beta}$, where $U_{\alpha}$ here is the domain of definition of $s_{\alpha}$ and $U_{\beta}$ is the domain of definition of $s_{\beta}$. The collection $\{g_{\alpha\beta}\}_{\alpha,\beta\in I}$ is termed the gluing 1-cocycle of the local frame field $\{s_{\alpha}\}_{\alpha\in I}$ of $\mathscr{L}$.
We then investigate the most basic properties of flat $\mathbb{R}$-line bundles.
@@
Let $M$ be a differentiable manifold and let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$. Then $\mathscr{L}\rightarrow M$ admits a local frame field $\{s_{\alpha}\}_{\alpha\in I}$ such that $Ds_{\alpha}=0$ for all $\alpha\in I$, and that the gluing 1-cocycle of the local frame field $\{s_{\alpha}\}_{\alpha\in I}$ consists of locally constant functions.
@@
Let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over a differentiable manifold $M$. A local frame field $\{s_{\alpha}\}_{\alpha\in I}$ of $\mathscr{L}\rightarrow M$ is termed an affine local frame field of $(\mathscr{L},D)$, if $Ds_{\alpha}=0$ for all $\alpha\in I$. The gluing 1-cocycle of some affine local frame field of $(\mathscr{L},D)$ is called a 1-cocycle of $(\mathscr{L},D)$.
@@
Two flat $\mathbb{R}$-line bundles $(\mathscr{L}_{1},D_{1})$ and $(\mathscr{L}_{2},D_{2})$ over a differentiable manifold $M$ are said to be gauge equivalent, if there exists a $\mathbb{R}$-vector bundle isomorphism $\phi\colon\mathscr{L}_{1}\rightarrow\mathscr{L}_{2}$ such that $D_{1}-\phi^{*}D_{2}\in\Omega^{1}(M)$ is an exact 1-form on $M$.
Definition 2.27.
Let $M$ be a differentiable manifold. The group of gauge equivalence classes of flat $\mathbb{R}$-line bundles over $M$ is termed the locally constant Picard group of $M$.
The following theorem, which relates the locally constant Picard group with cohomologies, is of fundamental importance.
Lemma 2.28.
Let $M$ be a differentiable manifold. The mapping that sends every flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ over $M$ to the Čech cohomology class of a 1-cocycle of $(\mathscr{L},D)$ in $\check{\mathrm{H}}(M;\mathbb{R}^{\times})$ is surjective, and two flat $\mathbb{R}$-line bundles over $M$ share the same image under this map if and only if they are gauge equivalent.
@@
This is a very special case of much more general statements in gauge theory. For a comprehensive proof, see [23]. ∎
We shall now have a closer look at the structure of the locally constant Picard group.
Proposition 2.29.
Let $M$ be a differentiable manifold. The induced homomorphism $w_{1}\colon\mathrm{H}(M;\mathbb{R}^{\times})\rightarrow\mathrm{H}(M;\mathbb{Z}/2\mathbb{Z})$ in the long exact sequence
$\cdots\longrightarrow\mathrm{H}^{1}(M;\mathbb{R})\longrightarrow\mathrm{H}^{1}(M;\mathbb{R}^{\times})\longrightarrow\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})\longrightarrow\cdots$
of cohomologies induced from the short exact sequence
$0\longrightarrow\mathbb{R}\overset{\operatorname{exp}}{\longrightarrow}\mathbb{R}^{\times}\overset{\operatorname{sgn}}{\longrightarrow}\mathbb{Z}/2\mathbb{Z}\longrightarrow 0$
of Abelian groups is precisely the first Stiefel-Whitney class.
Proof.
This statement follows from standard algebraic topology. ∎
A crucial difference between the exponential sequence in complex geometry and our short exact sequence
$0\longrightarrow\mathbb{R}\longrightarrow\mathbb{R}^{\times}\longrightarrow\mathbb{Z}/2\mathbb{Z}\longrightarrow 0$
is that the latter one splits. This allow us to define a new invariant for flat $\mathbb{R}$-line bundles.
Definition 2.30.
Let $M$ be a differentiable manifold, and let $\varphi\colon\mathbb{R}^{\times}\rightarrow\mathbb{R}$ be the even extension of the logarithm function, that is, $\varphi(x)=\operatorname{log}|x|$ for all $x\in\mathbb{R}^{\times}$. The induced homomorphism $\operatorname{sh}:=\varphi_{*}\colon\mathrm{H}(M;\mathbb{R}^{\times})\rightarrow\mathrm{H}(M;\mathbb{R})$ in the long exact sequence
$\cdots\longrightarrow\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})\longrightarrow\mathrm{H}^{1}(M;\mathbb{R}^{\times})\longrightarrow\mathrm{H}^{1}(M;\mathbb{R})\longrightarrow\cdots$
of cohomologies induced from the short exact sequence
$0\longrightarrow\mathbb{Z}/2\mathbb{Z}\longrightarrow\mathbb{R}^{\times}\overset{\operatorname{\varphi}}{\longrightarrow}\mathbb{R}\longrightarrow 0$
of Abelian groups is called the first Koszul class.
As an immediate consequence of Proposition 2.29 and Definition 2.30, the first Koszul class and the first Stiefel-Whitney class together classify the locally constant Picard group.
@@
Let $M$ be a differentiable manifold. The mapping
$w_{1}\oplus\operatorname{sh}\colon\mathrm{H}^{1}(M;\mathbb{R}^{\times})\rightarrow\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})\oplus\mathrm{H}^{1}(M;\mathbb{R}),\quad(\mathscr{L},D)\mapsto(w_{1}(\mathscr{L}),\operatorname{sh}(D))$
is an isomorphism of Abelian groups.
@@
Apply the splitting lemma and use the exactness of the cohomology functor. ∎
We finally arrived at the structure theorem of locally constant Picard groups.
Corollary 2.32.
Let $M$ be a differentiable manifold. The locally constant Picard group of $M$ is isomorphic to the additive group $\mathrm{H}^{1}(M;\mathbb{R})\oplus\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})$.
Proof.
The desired result follows from Lemma 2.28 and Corollary 2.31. ∎
In the case when the underlying manifold is aspherical, we may give an explicit description of the action of first Koszul class as an induced homomorphism.
Proposition 2.33.
Let $M$ be an aspherical differentiable manifold and $\pi_{1}(M)$ its fundamental group, so that we identify $\operatorname{Hom}(\pi_{1}(M),\mathbb{R})\equiv\mathrm{H}^{1}(M;\mathbb{R})$ via holonomy representation. Let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$, and let $\rho\colon\pi_{1}(M)\rightarrow\mathbb{R}^{\times}$ be its holonomy representation. Then, the homomorphism
$\pi_{1}(M)\rightarrow\mathbb{R},\quad g\mapsto\operatorname{log}|\rho(g)|$
as a group cohomology class in $\mathrm{H}^{1}(\pi_{1}(M);\mathbb{R})$ is precisely equal to the first Koszul class $\operatorname{sh}(D)$ of $(\mathscr{L},D)$.
Proof.
Denote by $G:=\pi_{1}(M)$ the fundamental group of $M$. Since $M$ is aspherical, the classifying map $M\to BG$ induces an isomorphism between the sheaf cohomology with locally constant coefficients and the corresponding group cohomology:
$\mathrm{H}^{1}(M;A)\cong\mathrm{H}^{1}(\pi_{1}(M);A)$
for any abelian group $A$.
Under this canonical isomorphism, the gauge equivalent class of $(\mathscr{L},D)$ in $\mathrm{H}^{1}(M;\mathbb{R}^{\times})$ corresponds to its holonomy representation $\rho\in\mathrm{H}^{1}(G;\mathbb{R}^{\times})\cong\operatorname{Hom}(G,\mathbb{R}^{\times})$.
Define group homomorphism $\varphi\colon\mathbb{R}^{\times}\rightarrow\mathbb{R}$ by $\varphi(x)=\operatorname{log}|x|$ for all $x\in\mathbb{R}^{\times}$. Then it readily seen that the following diagram commutes
$\begin{array}[]{ccc}\mathrm{H}^{1}(M;\mathbb{R}^{\times})&\overset{\operatorname{sh}}{\longrightarrow}&\mathrm{H}^{1}(M;\mathbb{R})\\ \downarrow&&\downarrow\\ \mathrm{H}^{1}(G;\mathbb{R}^{\times})&\overset{\varphi_{*}}{\longrightarrow}&\mathrm{H}^{1}(G;\mathbb{R})\end{array}$
and the desired property follows. ∎
We shall also be interested in a trivially criterion of elements of the locally constant Picard group. Recall that a line bundle is trivial if and only if it admits a nowhere vanishing global section. The following proposition gives a refinement in the case of flat $\mathbb{R}$-line bundles.
@@
To illustrate the power of the structure theorem of locally constant Picard group, we compare two proofs of the following torsion-free property.
Proposition 2.35.
Proof.
Proof.
Since $\mathscr{L}$ is a trivial $\mathbb{R}$-line bundle, its Stiefel-Whitney class $w_{1}(\mathscr{L})=0$. Therefore the gauge equivalent class of $(\mathscr{L},D)$ is contained in the subgroup $\mathrm{H}^{1}(M;\mathbb{R})$ of the locally constant Picard group of $M$. Since $(\mathscr{L}^{\otimes k},D)$ is trivial and $\mathrm{H}^{1}(M;\mathbb{R})$ is torsion-free, we conclude that $(\mathscr{L},D)$ is also trivial. ∎
2.3.2 Applications of Koszul Forms